OpenStudy (anonymous):

The area of a rectangle must be at least 78 square feet. If the length is 6 feet, find the minimum for the rectangle's width.

OpenStudy (anonymous):

area = 6 * w 78 = 6* w w = 78/6 = 13 feet

OpenStudy (anonymous):

thanks!! :)

OpenStudy (anonymous):

yw - plz click on good answer to close the question

OpenStudy (anonymous):

good answer box is to the right og my name - ita a rectangle with blue background